OpenStudy (anonymous):

What is the value of x when AB and CD are equidistant from the center of the circle and AB=3x+7 and CD=27-x

OpenStudy (anonymous):

equidistant chords are equal in size so AB = CD 3x + 7 = 27-x 3x+ x = 27 - 7 4x = 20 x = 20/4 x = 5